# ArtA: Automating Design Space Exploration of Spin Qubit Architectures

### v1

@@ VI.1 Termination Condition (TC)
During each optimization cycle, a specific termination condition (TC) is checked to reduce further computational time based on the way ArtA is used. Since we already obtained the ESP values for all architectures first during the third question, the highest ESP is known and can work as a TC. Additionally, we observed in practice that most runs reach the highest ESP sooner than 40 minutes; hence, this can be another TC. More specifically, our analysis revealed that over 95% of processes are completed earlier, while a minority extends significantly longer. Given that our evaluation focuses on runtime, omitting these prolonged instances won’t affect the overall assessment, as their inclusion would lead to unfavorable evaluations regardless.
@@ VI.2 Metrics Evaluating ArtA
To determine the performance of ArtA in the context of replying to the third research question, we need to define some relevant metrics. The optimization methods that will eventually determine ArtA’s performance should return a desired solution faster compared to searching with the brute-force approach. Otherwise, ArtA will be of no practical interest. We thus use the relative time-to-solution as a metric in Equation 6. The numerator includes the total runtime of both the optimization method and compilation until the desired result is reached (or the TC is triggered), and the denominator refers to the total runtime to brute force the entire design space.
@@
When the average compilation time increases due to larger quantum circuits, the execution time of the optimization method slowly becomes less important within the total runtime. Because of that, time-to-solution is not sufficient in evaluating ArtA’s performance. Thus, a relative calls-to-solution metric, in Eq. 7, is also a strong performance predictor for ArtA, especially for larger circuits. With calls-to-solution in the numerator, we refer to the total number of times SpinQ is called to compile for each new architecture. In case the optimization method attempts the same architecture twice or more, the compilation will be skipped since the results are already stored from the first call. Finally, the denominator represents the brute force approach and is the total number of architectures in the design space, which essentially equals $29,312$.
@@ VII Results and Evaluation
In this section, we will attempt to completely answer the third and fourth search questions. In both questions, we will focus on finding the architecture that obtains the highest ESP among all in the design space. However, before that, we briefly detail our experimental setup. Since all optimization methods are inherently probabilistic, multiple runs are required to create robust conclusions. Due to limited computational resources [^2] All time metrics were retrieved by running ArtA with Python 3.10.8 single-threaded on a 2017 MacBook Pro, using a 3.1 GHz Quad Core Intel i7 processor and 16 GB of RAM., we ran each method for each hyper-parameter combination ten times alongside the TC. Besides the mean performance, we will also report the worst performing results for reference.
@@ VII.1 Comparison of optimization methods
In this first part of our results, we will address the third research question stated in Section III, which is about finding the best performant optimization method according to our ArtA metrics defined in Section VI.2. The results of the relative time-to-solution metric for finding the architecture with the highest ESP are given in Table 3. This table also provides insights into matching different method configurations (described in Appendix X.2) with specific quantum circuits. As explained previously, the denominator of Eq. 6 is obtained beforehand by measuring the time taken by the brute force approach. We observe overall that for each quantum circuit, on average, all optimization methods find the solution in less time than brute-forcing (values under 100). For a better interpretation of the numbers on the table, a 50 means the particular optimization configuration of ArtA takes half the time of brute forcing all architecture, and 100 means that it takes the same time.
Multiple optimization methods greatly outperform the random sampling (RS column), as expected. The best performance is at 0.9% on average for QV10 with SA [20, 3], meaning that ArtA, with this optimization method configuration and quantum circuit, takes only 0.9% on average of the total brute force time to find the architecture with the highest ESP.
We calculated based on Table 3 that, on average, for all optimization methods, ArtA was able to decrease the exploration time by 80.6%. Overall, different optimization methods clearly show varied performance across their respective columns. However, the performance differences among various hyper-parameter settings within a single optimization method are less distinct. In a performance order, we can observe GA being first, followed by BO, ACO, and, lastly, SA. More specifically, GA, with a population size of 50 and a mutation probability of 0.15, finds the highest ESP the fastest at 11.9% relative time-to-solution, on average, with BO [EI, MA] coming close at 12.1%. Lastly, in 86.7% of the cases, a top 5 optimization method configuration in the worst row is also a top 5 optimization method configuration in the mean row, which indicates the consistency of the results. To further support our findings, next, we will assess the relative calls-to-solution metric.
@@
From the different variants of the GA and BO used, we focus on GA [100, 0.2] and BO [EI, MA], as they have the best performance and did not reach our TC in any of the ten runs. By comparing the two tables, we can observe that BO [EI, MA] needs marginally more time to evaluate fewer architectures compared to GA [100, 0.2]. Since the compilation time per architecture is the same between methods, this points to a slightly (0.2% on average) longer execution time of BO [EI, MA] compared to GA [100, 0.2]. This can be explained based on the algorithm’s reliance on matrix multiplication and several other linear algebra operations, which are computationally more expensive than the more simple operations of GA (see Sec. X.2). However, the actual total runtime of ArtA will depend on the ratio between compilation and optimization time, which can change for large circuits. Based on their small relative time-to-solution difference, we can predict that BO [EI, MA] will outperform GA [100, 0.2] for larger circuits as the latter has, on average, 58% higher relative calls-to-solution. Therefore, we conclude that BO [EI, MA] is a better candidate for automating the DSE process for large-scale circuits, as it can traverse the design space faster.
@@ VII.2 Architecture Analysis
In Table 5, we present the values of the architectural variables, introduced in Sec. IV, that were found by ArtA having the highest ESP with the BO [EI, MA] optimization method alongside their ESP values. These circuits are organized by class and size in number of qubits within the rows of the table. Subsequent columns present the ESP of the best architecture discovered, followed by color-coded architectural parameter categories as delineated in Sec. IV. The use of color accents in the table serves to enhance visual clarity, with each corresponding to the values of each variable.
@@
First, we observe a unanimous preference for the Z rotation implementation (z_rot_impl) in the zri column, with a pulse-based rotation being preferred. Note that previously, we have assumed in Section V.1 that a shuttle and a pulse-based Z rotation share the same single-qubit gate fidelity, but because the latter requires two shuttles, it can benefit the ESP more than the other type. Moreover, as a shuttle operation involves two quantum dots while a pulse-based operation only involves one, the shuttle-based Z rotation will have a higher chance of incurring crosstalk errors based on our crosstalk definition. Given these two reasons, BO has favored the shuttle-based version. This indicates that shuttle-based single-qubit rotations are not as sustainable as pulsed-based for large-scale circuits.
Moving on, the X, Y, and Z gates are implemented the same way hardware-wise, xy_z automatically is assumed 1 based on the interdependencies detailed in Section IV.6. Consequently, xy_tqg and z_tqg are equivalent, taking both either 0 or 1 from the optimization method of ArtA (BO [EI, MA]), as we can observe in the table. Additionally, xy_z_tqg can also be conceptually equivalent to xy_tqg and z_tqg since the pulse-based Z rotation is selected. Therefore, whenever xy_z_tqg is 1 or xy_tqg and z_tqg is 1, it means the architecture(s) which allow(s) parallelization of all single- and two-qubit gates obtains the highest ESP. In that sense, one would expect xy_z_tqg, xy_tqg and z_tqg to be synchronized. However, since such exploration freedom was given in the design space, the reasons for the mismatches could be explained by the internal characteristics of the circuits themselves. For example, in QFT100, the parallelization of all gate types occurs more often than in QFT60, which has xy_z_tqg = 0, xy_tqg = 1, and z_tqg = 1.
Conversely, when applying the same reasoning to QV8, QV10, and QFT20, it is not necessary for the hardware to support parallel executions of single-qubit and two-qubit gates. This shows that the success of real quantum circuits can come from different “angles” with reduced hardware capabilities  [^3] Ultimately, each circuit profile has specific architectural traits that can be leveraged to achieve optimal performance without the need for excessively advanced hardware features (achieving a balance in architectural characteristics)..
@@ IX Conclusion
In this paper, we propose a comprehensive exploration of the current and future architectural design space for spin-qubit-based quantum dots quantum processors. Our research focuses on assessing the critical architectural characteristics that could be key in ensuring high performance on such devices. To this end, we present an upgraded version of the SpinQ compilation framework for spin qubit architectures in which architectural variables can configure internal compilation passes, making it the first Design Space Exploration (DSE) framework. After defining our design space, consisting of $29,312$ architectures, we propose a multi-optimization-based tool, ArtA (Artificial Architect), to automate the DSE process. The goal of ArtA is (a) to suggest which of the seventeen optimization method configurations can find the architecture with the highest ESP the fastest compared to the brute forcing approach and (b) which architectural design characteristics are key for building high-performance spin-qubit devices per circuit category. We have shown up to 99.1% improvement in computation times compared to the brute-force approach. We have also provided insights into matching different optimization methods with specific quantum circuit categories. Bayesian optimization with an Expected Improvement acquisition function and a Matérn kernel was chosen as the best optimization method due to its effectiveness in managing higher qubit counts. After that, we proceed to automate the DSE process for 44 quantum circuits of up to 258 qubits within the 29,312-architecture design space. Our findings reveal the importance of minimizing the circuit duration by maximizing parallelization whenever possible instead of selecting architectural characteristics that minimize crosstalk errors. Crosstalk is currently a significant concern in experiments, but our long-term considerations deem decoherence more crucial. Lastly, the shuttle operation is preferred over costly SWAP operations for communication purposes, especially for large-scale circuits. Contrary to that, shuttle-based single-qubit gates such as Z rotations are not preferred over pulsed-based. Furthermore, our universal architecture suggestion places higher priority on parallelizing single-qubit gates rather than two-qubit gates.
@@
As for future improvements, a broader range of hyper-parameter combinations and circuit variations could enhance ArtA’s effectiveness. Then, with advancements in spin qubit device fabrication, new architectural variables, and values could potentially update the design space, and such integration will be relatively straightforward owing to the modularity of the internal functions of SpinQ and ArtA. Moreover, introducing non-architectural constraints such as development costs and time into ArtA’s design space could help make real-world conclusions regarding developing spin qubits. This would allow for operational fidelities to become variables as well and, consequently, add a second dimension to the Figure of Merit (ESP).

### v2

@@ VI.1 Termination Condition (TC)
During each optimization cycle, a specific termination condition (TC) is checked to reduce further computational time based on the way ArtA is used. Since we first obtain the ESP values for all architectures during the third question, the highest ESP is known and can work as a TC. Additionally, we observed in practice that most runs reach the highest ESP sooner than 40 minutes; hence, this can be another TC. More specifically, our analysis revealed that over 95% of processes are completed earlier, while a minority extends significantly longer. Given that our evaluation focuses on runtime, omitting these prolonged instances won’t affect the overall assessment, as their inclusion would lead to unfavorable evaluations regardless.
@@ VI.2 Metrics Evaluating ArtA
To determine the performance of ArtA in the context of replying to the third research question, we need to define some relevant metrics. The optimization methods that will eventually determine ArtA’s performance should return a desired solution faster compared to searching with the brute-force approach of SpinQ. Otherwise, ArtA will be of no practical interest. We thus use the relative time-to-solution as a metric in Equation 6. The numerator includes the total runtime of both the optimization method and compilation until the desired result is reached (or the TC is triggered), and the denominator refers to the total runtime to brute force the entire design space.
@@
When the average compilation time increases due to larger quantum circuits, the execution time of the optimization method slowly becomes less important within the total runtime. Because of that, time-to-solution is not sufficient in evaluating ArtA’s performance. Thus, a relative calls-to-solution metric, in Equation 7, is also a strong performance predictor for ArtA, especially for larger circuits. With calls-to-solution in the numerator, we refer to the total number of times SpinQ is called to compile for each new architecture. In case the optimization method attempts the same architecture twice or more, the compilation will be skipped since the results are already stored from the first call. Finally, the denominator represents the brute force approach and is the total number of architectures in the design space, which essentially equals $29,312$.
@@ VII Results and Evaluation
In this section, we will attempt to completely answer the third and fourth research questions. In both questions, we will focus on finding the architecture that obtains the highest ESP among all in the design space. Since all optimization methods are inherently probabilistic, multiple runs are required to create robust conclusions. Due to limited computational resources [^2] All time metrics were retrieved by running ArtA with Python 3.10.8 single-threaded on a 2017 MacBook Pro, using a 3.1 GHz Quad Core Intel i7 processor and 16 GB of RAM., we ran each method for each hyper-parameter combination ten times alongside the TC. Besides the mean performance, we will also report the worst performing results for reference.
@@ VII.1 Comparison of optimization methods
In this first part of our results, we will address the third research question stated in Section III, which is about finding the best performant optimization method according to our ArtA metrics defined in Section VI.2. The results of the relative time-to-solution metric for finding the architecture with the highest ESP are given in Table 3. This table also provides insights into matching different method configurations (described in Appendix X.2) with specific quantum circuits. As explained previously, the denominator of Equation 6 is obtained beforehand by measuring the time taken by the brute force approach. We observe overall that for each quantum circuit, on average, all optimization methods find the solution in less time than brute-forcing (values under 100). For a better interpretation of the numbers on the table, a 50 means the particular optimization configuration of ArtA takes half the time of brute forcing all architecture, and 100 means that it takes the same time.
Multiple optimization methods greatly outperform the random sampling (RS column), as expected. The best performance is at 0.9% on average for QV10 with SA [20, 3], meaning that ArtA, with this optimization method configuration and quantum circuit, takes only 0.9% relative time-to-solution, on average, of the total brute force time to find the architecture with the highest ESP.
We calculated based on Table 3 that, on average, for all optimization methods, ArtA was able to decrease the exploration time by 80.6%. Overall, different optimization methods clearly show varied performance across their respective columns. However, the performance differences among various hyper-parameter settings within a single optimization method are less distinct. In a performance order, we can observe GA being first, followed by BO, ACO, and, lastly, SA. More specifically, GA, with a population size of 50 and a mutation probability of 0.15, finds the highest ESP the fastest at 11.9% relative time-to-solution, on average, with BO [EI, MA] coming close at 12.1%. Lastly, in 86.7% of the cases, a top 5 optimization method configuration in the worst row is also a top 5 optimization method configuration in the mean row, which indicates the consistency of the results across the ten runs. To further support our findings, next, we will assess the relative calls-to-solution metric.
@@
From the different variants of the GA and BO used, we focus on GA [100, 0.2] and BO [EI, MA], as they have the best performance and did not reach our TC in any of the ten runs. By comparing the two tables, we can observe that BO [EI, MA] needs marginally more time to evaluate fewer architectures compared to GA [100, 0.2]. Since the compilation time per architecture is the same between methods, this points to a slightly (0.2% on average) longer execution time of BO [EI, MA] compared to GA [100, 0.2]. This can be explained based on the algorithm’s reliance on matrix multiplication and several other linear algebra operations, which are computationally more expensive than the more simple operations of GA (see Section X.2). However, the actual total runtime of ArtA will depend on the ratio between compilation and optimization time, which can change for large circuits. Based on their small relative time-to-solution difference, we can predict that BO [EI, MA] will outperform GA [100, 0.2] for larger circuits as the latter has, on average, 58% higher relative calls-to-solution. Therefore, we conclude that BO [EI, MA] is a better candidate for automating the DSE process for large-scale circuits, as it can traverse the design space faster.
@@ VII.2 Architecture Analysis
In Table 5, we present the values of the architectural variables, introduced in Section IV, that were found by ArtA having the highest ESP with the BO [EI, MA] optimization method alongside their ESP values. These circuits are organized by class and size in number of qubits within the rows of the table. Subsequent columns present the ESP of the best architecture discovered, followed by color-coded architectural parameter categories as delineated in Section IV. The use of color accents in the table serves to enhance visual clarity, with each corresponding to the values of each variable.
@@
First, we observe a unanimous preference for the Z rotation implementation (z_rot_impl) in the zri column, with a pulse-based rotation being preferred. Note that previously, we have assumed in Section V.1 that a shuttle and a pulse-based Z rotation share the same single-qubit gate fidelity, but because the latter requires two shuttles, it can benefit the ESP more than the other type. Moreover, as a shuttle operation involves two quantum dots while a pulse-based operation only involves one, the shuttle-based Z rotation will have a higher chance of incurring crosstalk errors based on our crosstalk definition. Given these two reasons, BO has favored the pulsed-based version. This indicates that shuttle-based single-qubit rotations are not as sustainable as pulsed-based for large-scale circuits.
Moving on, since the X, Y, and Z gates are implemented the same way hardware-wise, xy_z automatically is assumed 1 based on the interdependencies detailed in Section IV.6. Consequently, xy_tqg and z_tqg are equivalent, taking both either 0 or 1 from the optimization method of ArtA (BO [EI, MA]), as we can observe in the table. Additionally, xy_z_tqg can also be conceptually equivalent to xy_tqg and z_tqg since the pulse-based Z rotation is selected. Therefore, whenever xy_z_tqg is 1 or xy_tqg and z_tqg is 1, it means the architecture(s) which allow(s) parallelization of all single- and two-qubit gates obtains the highest ESP. In that sense, one would expect xy_z_tqg, xy_tqg and z_tqg to be synchronized. However, since such exploration freedom was given in the design space, the reasons for the mismatches could be explained by the internal characteristics of the circuits themselves. For example, in QFT100, the parallelization of all gate types occurs more often than in QFT60, which has xy_z_tqg = 0, xy_tqg = 1, and z_tqg = 1.
Conversely, when applying the same reasoning to QV8, QV10, and QFT20, it seems unnecessary for the hardware to support parallel executions of single-qubit and two-qubit gates. This shows that the success of real quantum circuits can come from different “angles” with reduced hardware capabilities  [^3] Ultimately, each circuit profile has specific architectural traits that can be leveraged to achieve optimal performance without the need for excessively advanced hardware features (achieving a balance in architectural characteristics)..
@@ IX Conclusion
In this paper, we propose a comprehensive exploration of the current and future architectural design space for quantum dot spin-qubit quantum processors. Our research focuses on assessing the critical architectural characteristics that could be key in ensuring high performance on such devices. To this end, we present an upgraded version of the SpinQ compilation framework for spin qubit architectures in which architectural variables can configure internal compilation passes, making it the first Design Space Exploration (DSE) framework. After defining our design space, consisting of $29,312$ architectures, we propose a multi-optimization-based tool, ArtA (Artificial Architect), to automate the DSE process. The goal of ArtA is (a) to suggest which of the seventeen optimization method configurations can find the architecture with the highest ESP the fastest compared to the brute forcing approach and (b) which architectural design characteristics are key for building high-performance spin-qubit devices per circuit category. We have shown up to 99.1% improvement in computation times compared to the brute-force approach, showing that it is possible to easily explore a vast number of designs. We have also provided insights into matching different optimization methods with specific quantum circuit categories. Bayesian optimization with an Expected Improvement acquisition function and a Matérn kernel was chosen as the best optimization method due to its effectiveness in managing higher qubit counts. After that, we proceed to automate the DSE process for 44 quantum circuits of up to 258 qubits within the 29,312-architecture design space. Our findings reveal the importance of minimizing the circuit duration by maximizing parallelization whenever possible instead of selecting architectural characteristics that minimize crosstalk errors. Crosstalk is currently a significant concern in experiments, but our long-term considerations deem decoherence more crucial. Lastly, the shuttle operation is preferred over costly SWAP operations for communication purposes, especially for large-scale circuits. Contrary to that, shuttle-based single-qubit gates such as Z rotations are not preferred over pulsed-based. Furthermore, our universal architecture suggestion places higher priority on parallelizing single-qubit gates rather than two-qubit gates. Overall, our work illustrates how the integration of DSE methodologies with optimization algorithms can be successfully utilized to offer valuable insights to designers of quantum processors.
@@
As for future improvements, a broader range of hyper-parameter combinations and circuit variations could enhance ArtA’s effectiveness. Then, with advancements in spin qubit device fabrication, new architectural variables and values could potentially update the design space, and such integration will be relatively straightforward owing to the modularity of the internal functions of SpinQ and ArtA. Moreover, introducing non-architectural constraints such as development costs and time into ArtA’s design space could help make real-world conclusions regarding developing spin qubits. This would allow for operational fidelities to become variables as well and, consequently, add a second dimension to the Figure of Merit (ESP).